OpenStudy (anonymous):

f(s) = (√s - 1) / (√s + 1) f'(s) = ?

OpenStudy (anonymous):

It'll be easier to change the square root to s^(1/2)... then use the quotient rule.

OpenStudy (anonymous):

ugly algebra involved for sure \[(\frac{f}{g})'=\frac{gf'-fg'}{g^2}\] use \[f(x)=\sqrt{x-1},f'(x)=\frac{1}{2\sqrt{x-1}}\] \[g(x)=\sqrt{x+1},g'(x)=\frac{1}{2\sqrt{x+1}}\] and make the replacements

OpenStudy (anonymous):

btw it is good to remember that \[\frac{d}{dx}\sqrt{x}=\frac{1}{2\sqrt{x}}\] it comes up frequently and it is always the same, sort of like memorizing 7 times 8

OpenStudy (anonymous):

Sometimes it's easier to find the derivative by making the question look a little simpler for yourself. If it makes it easier and you're more comfortable with the expansion with the multiplying rule, instead of seeing \[\sqrt{s}\] think of it as \[s^{1/2}\] Make it \[f(s) = (s^{1/2} -1) (s ^{1/2} +1)^{-1}\]

OpenStudy (anonymous):

so what is s^1/2 times 1/2 s^-1/2 ??

OpenStudy (anonymous):

i completely disagree

OpenStudy (anonymous):

f(s)=(s1/2−1)(s1/2+1)−1 would mean i would have to use the chain rule, and i want to solve this using the quotient rule

OpenStudy (anonymous):

Then you would follow what satellite used

OpenStudy (anonymous):

also i am assuming you meant \[\frac{\sqrt{x-1}}{\sqrt{x+1}}\] although i could be wrong

OpenStudy (anonymous):

*Chain rule, the name slipped my mind for a second

OpenStudy (anonymous):

no, only the sq. roots of x in each of those

OpenStudy (anonymous):

oooh in that case it is much easier

OpenStudy (anonymous):

\[f(x)=\sqrt{x}-1,f'(x)=\frac{1}{2\sqrt{x}}\] \[g(x)=\sqrt{x}+1,g'(x)=\frac{1}{2\sqrt{x}}\]

OpenStudy (anonymous):

\[\frac{(\sqrt{x}+1)\times \frac{1}{2\sqrt{x}}-(\sqrt{x}-1)\times \frac{1}{2\sqrt{x}}}{(\sqrt{x}+1)^2}\]

OpenStudy (valpey):

\[f(s) = \frac{s^{1/2}-1}{s^{1/2}+1}\] \[f'(s) = \frac{1}{(s^{1/2}+1)^2s^{1/2}}\]

OpenStudy (anonymous):

Alright so here is the answer using only the division rule: u = s^1/2 - 1 u' = (1/2)s^-(1/2) v = s^1/2 + 1 v' = (1/2)s^-(1/2) f'(s) = (s^1/2 + 1)((1/2)s^-(1/2)) - (s^1/2 - 1)((1/2)s^-(1/2)) / (√s + 1)² = (1/2) + (1/2s^-(1/2)) - ((1/2) - ((1/2s^-(1/2) / (√s + 1)² = (1/2s^-(1/2) + (1/2s^-(1/2) / (√s + 1)² = s^-(1/2) / (√s + 1)² = 1/ √s(√s+1)²

OpenStudy (anonymous):

multiply top and bottom by \[\frac{1}{2\sqrt{x}}\] to clear the fractions and get \[\frac{\sqrt{x}+1-(\sqrt{x}-1)}{2\sqrt{x}(\sqrt{x}+1)^2}\] \[=\frac{2}{2\sqrt{x}(\sqrt{x}+1)^2}\] \[=\frac{1}{\sqrt{x}(\sqrt{x}+1)^2}\]

OpenStudy (anonymous):

you got it!

OpenStudy (anonymous):

there is in fact a somewhat snappier way to do it using the chain rule, but it might not be worth the effort

OpenStudy (anonymous):

yes it took me a long time to realize that when you multiply s^1/2 by 1/2s^-(1/2) that the answer is just 1/2 because you add the exponents together so (1/2) + -(1/2) = 0 and then s^0 is just 1 so s^1/2 * 1/2s^-(1/2) = 1/2

OpenStudy (anonymous):

yes it is true. it might be easier to do this problem \[\frac{d}{dx}\frac{x-1}{x+1}\] and get \[\frac{2}{(x+1)^2}\] then replace x by \[\sqrt{x}\] and multiply by \[\frac{1}{2\sqrt{x}}\] but that requires some thinking which might not be clear

OpenStudy (anonymous):

The hardest part of this question was knowing where to start and how to rearrange the original equation to work step by step in order to simplify things